# Probabilistic Kernel Function for Fast Angle Testing

### v1

@@ 4.2 A New Probabilistic Test in Similarity Graph
This is the original task of PEOs peos , where the authors introduce probabilistic routing into similarity graphs to accelerate the search. The definition of probabilistic routing is as follows.
Definition 4.1 (Probabilistic Routing peos )
Given a query vector $\bm{q}$, a node $v$ in the graph index, an error bound $\epsilon$, and a distance threshold $\delta$, for an arbitrary neighbor $w$ of $v$ such that $dist(\bm{w},\bm{q})<\delta$, if a routing algorithm returns true for $w$ with a probability of at least $1-\epsilon$, then the algorithm is deemed to be $(\delta,1-\epsilon)$-routing.
In peos , the authors proposed a $(\delta,1-\epsilon)$-routing test called PEOs test. Here, based on $K^{2}_{S}$, we propose a new routing test for $\ell_{2}$ distance, called the KS2 test, as follows.
@@
Here, $\bm{q}\in\mathbb{R}^{d}$ is the query, $v$ is the visited graph node, $w$ is the neighbor of $v$ and $e$ is the edge between $v$ and $w$. $\tau$ is the threshold determined by the result list of graph. $\bm{q_{i}}$, $\bm{e_{i}}$, $\bm{u^{i}_{j}}$ denote the $i$-th sub-vectors of $\bm{q}$ ($1\leq i\leq L$), $\bm{e}$ and $\bm{u_{j}}$, respectively. $\bm{u^{i}_{e[i]}}$ denotes the reference vector of $\bm{e_{i}}$ among all $\{\bm{u^{i}_{j}}\}$’s ($1\leq j\leq m$). In our experiments, $S$ is taken to be $S_{\text{sym}}(256,L)$.
During the traversal of the similarity graph, we check the exact distance from graph node $w$ to $q$ only when ineq. (10) is satisfied; otherwise, we skip the computation of $w$ for efficiency. A complete graph-based algorithm equipped with the KS2 test can be found in Alg. 6. By Lemma 3.3 and the same analysis in peos , we can easily obtain the following result.
Corollary 4.2
@@
Complexity analysis. For the time complexity, for every edge $e$, the computation of the LHS of ineq. (10) requires $L$ lookups in the table and $L-1$ additions, while the computation of the RHS of ineq. (10) requires two additions and one multiplication. By using SIMD, we can perform the KS2 test for $16$ edges simultaneously. For the space complexity, for each edge, we need to store $L$ bytes to recover $\bm{{q_{i}}^{\top}u^{i}_{e[i]}}$, along with two scalars, that is, $A_{S}(\bm{v})\|\bm{w}\|^{2}/(2\|\bm{e}\|)$ and $A_{S}(\bm{v})/\|\bm{e}\|$, which are quantized using scalar quantization to enable fast computation of the RHS of ineq. (10).
@@ 5 Experiments
All experiments were conducted on a PC equipped with an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz. KS1 and KS2 were implemented in C++. The ANNS experiments used 64 threads for indexing and a single CPU for searching. We evaluated our methods on six high-dimensional real-world datasets: Word, GloVe1M, GloVe2M, Tiny, GIST, and SIFT. Detailed statistics for these datasets are provided in Appendix C.1. More experimental results can be found in Appendix C.
@@ 5.1 Comparison with CEOs
As demonstrated in Sec. 4.1, the results of CEOs are directly used to accelerate other similarity search processes Falconn++ ; sDBSCAN . In this context, we focus solely on the improvement of CEOs itself. Specifically, we show that KS1, equipped with the structures $S_{\text{sym}}(m,1)$ and $S_{\text{pol}}(m,1)$, can slightly outperform CEOs($m$) on the original task of CEOs, that is, $k$-MIPS, where $m$ denotes the number of projection vectors and was set to 2048, following the standard configuration of the original CEOs. Since the only difference among the compared approaches is the configuration of the projection vectors, we use a unified algorithm (see construction of projection structure Alg. 4 and MIPS query processing Alg. 5 in Appendix) with the configuration of projection vectors as an input to compare their recall rates. From the results in Tab. 1, we observe that: (1) in most cases, KS1 with the two proposed structures achieves slightly better performance than CEOs, supporting our claim that a smaller reference angle yields a more accurate estimation, and (2) $S_{\text{pol}}$ generally achieves a higher recall rate than $S_{\text{sym}}$, verifying that a configuration closer to the best covering yields better performance.
Table 1: Comparison of recall rates (%) for $k$-MIPS, $k$ = 10. The number of projection vectors was fixed at 2048 for all compared methods. Top-5 projection vectors are probed. Probe@$n$ indicates that the top-$n$ points were probed on each probed projection vector. To eliminate the bias introduced by random projection, we obtain the results over 10 runs and report the average recall rate.
@@ 5.2 ANNS Performance
We chose ScaNN Scann , HNSW hnsw , and HNSW+PEOs peos as our baselines, where ScaNN is a state-of-the-art quantization-based approach that performs better than IVFPQFS, and HNSW+PEOs peos is a state-of-the-art graph-based approach that outperforms FINGER FINGER and Glass. Similar to HNSW+PEOs, KS2 is implemented on HNSW, and the corresponding approach is named HNSW+KS2. The detailed parameter settings of all compared approaches can be found in Appendix C.1, and additional experimental results can also be found in Appendix C.
@@
Figure 2: Impact of $L$ (See Appendix C.4 for other datasets). $k$ = 10. The y-axis of the upper figures denotes the additional index cost (%) of HNSW+PEOs compared to the original HNSW.
@@ 6 Conclusions
Appendix A Proof of Lemmas
A.1 Proof of Lemma 3.2
@@ Appendix A Proof of Lemmas
$\mathbb{P}[K^{2}_{S}(\bm{q},\bm{v})\geq\cos\theta]=I_{t^{\prime}}(\frac{d-2}{2},\frac{d-2}{2})$ (25)
@@
Instead of $\bm{v}$ and $\bm{u}$, we consider $-\bm{v}$ and $-\bm{u}$, and construct the simplex based on $\bm{q}$, $\bm{-v}$ and $\bm{-u}$ as in Case 1. Then, with the reverse of sign, we finally obtain an equation similar to eq. (22), except with $\alpha$ replaced by $\pi-\alpha$. The following analysis is similar to that of Case 1.
@@
A.3 Proof of Lemma 3.4
Before proving this theorem, we first introduce a uniformly distributed sequence.
Definition A.1
(Uniformly distributed sequence) A sequence $\{w_{N}\}^{\infty}_{N=2}$ of $N$-point configuration on $\mathbb{S}^{d-1}$ is called uniformly distributed if for every closed subset $B\in\mathbb{S}^{d-1}$ whose boundary relative to $\mathbb{S}^{d-1}$ has $d$-dimensional measure zero,
$\lim_{N\to\infty}\frac{\#(\omega_{N}\cap B)}{N}=\sigma_{d}(B)$ (30)
where $\sigma_{d}(B)$ is the area measure on $\mathbb{S}^{d-1}$ normalized to a probability measure.
Proof Let $\{v_{i,N}\}^{N}_{i=1}\subset\mathbb{S}^{d-1}$ be an i.i.d. sequence drawn from the uniform distribution $\sigma_{d}$, i.e., the area measure on $\mathbb{S}^{d-1}$ normalized to be a probability measure, and let $w_{N}=\{v_{1,N},v_{2,N}\ldots v_{N,N}\}$.
By SLLN, we have
[…]
$\mathbb{P}\left[\lim_{N\to\infty}\frac{\#(\omega_{N}\cap B)}{N}=\sigma_{d}(B)\right]=1.$ (32)
On the other hand, the sequence $\{w_{N}\}^{\infty}_{N=2}$ is uniformly distributed on $\mathbb{S}^{d-1}$ if and only if, for every continuous function $f:\mathbb{S}^{d-1}\rightarrow\mathbb{R}$, we have
$\frac{1}{N}\sum_{i=1}^{N}f(v_{i,N})\rightarrow\int_{\mathbb{S}^{d-1}}f(v)\,\mathrm{d}\sigma_{d}(v),\quad N\rightarrow\infty.$ (33)
Define $f(v)$ as follows
$f(v)=\cos\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle.$ (34)
Clearly, $f(v)$ is a continuous function w.r.t $v$. if we substitute $f(v)$ into relationship (33), the lemma is proved.
@@
A.4 Proof of Lemma 3.5
@@
Then we only need to focus on $J(S_{\text{ran}}(m,L))$ and prove that it is equal to the RHS of ineq. (8). Let $\bm{v=[v_{1},v_{2},\cdots.v_{L}]}\in\mathbb{R}^{n}$ be a vector drawn randomly from $\mathbb{S}^{d-1}$, where $d$ is assumed to be divisible by $L$ and $d=Ld^{\prime}$. Let $r_{L}(\bm{v})$ be the regularized vector w.r.t. $\bm{v}$. That is,
@@
Next, we consider $\mathbb{E}[T^{\prime}(d^{\prime},L)]$. Because of the rotational symmetry of the sphere, the distribution of the inner product $Z=\langle u,v\rangle$ (for fixed $v$ and uniformly random $u$) depends only on the dimension $d^{\prime}$. It has the following density on $[-1,1]$:
@@
Appendix B Supplementary Materials
B.1 Illustration of Random Projection Techniques for Angle Estimation
@@
B.2 ANNS and Similarity Graphs
As one of the most fundamental problems, approximate nearest neighbor search (ANNS) has seen a surge of interest in recent years, leading to the development of numerous approaches across different paradigms. These include tree-based methods [10], hashing-based methods [3, 20, 1, 4, 26], vector quantization (VQ)-based methods [19, 15, 5, 16], learning-based methods [17, 21], and graph-based methods [23, 27, 14, 13].
Among these, graph-based methods are widely regarded as the state-of-the-art. Currently, three main types of optimizations are used to enhance their search performance: (1) improved edge-selection strategies, (2) more effective routing techniques, and (3) quantization of raw vectors. These approaches are generally orthogonal to one another. Since the proposed structure, PEOs, belongs to the second category, we briefly introduce several highly relevant works below. TOGG-KMC[29] and HCNNG[24] use KD-trees to determine the direction of the query and restrict the search to points in that direction. While this estimation is computationally efficient, it results in relatively low query accuracy, limiting their improvements over HNSW[23]. FINGER[9] examines all neighbors and estimates their distances to the query. Specifically, for each node, FINGER locally generates promising projection vectors to define a subspace, then uses collision counting which is similar to SimHash to approximate distances within each visited subspace. Learn-to-Route [6] learns a routing function using auxiliary representations that guide optimal navigation from the starting vertex to the nearest neighbor. Recently, the authors of [22] proposed PEOs, which leverages space partitioning and random projection techniques to estimate a random variable representing the angle between each neighbor and the query vector. By aggregating projection information from multiple subspaces, PEOs substantially reduces the variance of this estimated distribution, significantly improving query accuracy.
B.3 Algorithms Related to KS1 and KS2

### v2

@@ 6.2 A New Probabilistic Test in Similarity Graph
Lu et al. (2024) proposed probabilistic routing and used it in similarity graphs to accelerate ANNS. Let $dist(\cdot,\cdot)$ be the distance function for ANNS. Each node in the similarity graph represents a data vector. The definition of probabilistic routing is as follows.
Definition 6.1 (Probabilistic Routing (Lu et al., 2024)).
Given a query vector $\bm{q}$, a node $v$ in the graph, an error bound $\epsilon$, and a distance threshold $\delta$, for an arbitrary neighbor $w$ of $v$ such that $dist(\bm{w},\bm{q})<\delta$, if a routing algorithm returns true for $w$ with a probability of at least $1-\epsilon$, then the algorithm is deemed to be $(\delta,1-\epsilon)$-routing.
Lu et al. (2024) proposed a $(\delta,1-\epsilon)$-routing test called PEOs test. Based on $K^{2}_{S}$, we propose a new routing test for $\ell_{2}$ distance, called the KS2 test, as follows (see Sec. C.1 in Appendix for more detail).
@@
$\bm{q}\in\mathbb{R}^{d}$ is the query, $\bm{v}$ is the visited graph node, $\bm{w}$ is the neighbor of $\bm{v}$, and $\bm{e}=\bm{w}-\bm{v}$. $\tau$ is the threshold determined by the result list of graph. $\bm{q_{i}}$, $\bm{e_{i}}$ denote the $i$-th sub-vectors of $\bm{q}$ and $\bm{e}$, respectively ($1\leq i\leq L$). $\bm{u}^{i}_{j}$ denotes the $j$-th element of the $i$-th sub-vector of $\bm{u}$. $\bm{u^{i}_{e[i]}}$ denotes the reference vector of $\bm{e_{i}}$ among all $\{\bm{u^{i}_{j}}\}$’s ($1\leq j\leq m$). In our experiments, $S$ was set to $S_{\text{sym}}(256,L)$.
During the traversal of the similarity graph, we check the exact distance from graph node $w$ to $q$ only when Ineq. (9) is satisfied; otherwise, we skip the computation of $w$ for efficiency. A complete graph-based algorithm equipped with the KS2 test can be found in Alg. 6. By Lemma 4.3 and the same analysis in Lu et al. (2024), we can easily obtain the following result.
Corollary 6.2.
@@
Complexity analysis. For the time complexity, for every edge $\bm{e}$, the computation of the LHS of Ineq. (9) requires $L$ lookups in the table and L1L-1 additions, while the computation of the RHS of Ineq. (9) requires two additions and one multiplication. By using SIMD, we can perform the KS2 test for $16$ edges simultaneously. For the space complexity, for each edge, we need to store $L$ bytes to recover $\bm{{q_{i}}^{\top}u^{i}_{e[i]}}$, along with two scalars, that is, $A_{S}(\bm{v})\|\bm{w}\|^{2}/(2\|\bm{e}\|)$ and $A_{S}(\bm{v})/\|\bm{e}\|$, which are quantized using scalar quantization to enable fast computation of the RHS of Ineq. (9).
@@ 7 Experiments
All experiments were conducted on a PC equipped with an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz. KS1 and KS2 were implemented in C++. The ANNS experiments used 64 threads for indexing and a single CPU for searching. We evaluated our methods on six high-dimensional real-world datasets: Word, GloVe1M, GloVe2M, Tiny, GIST, and SIFT. Detailed statistics for these datasets are provided in Appendix D.1. More experimental results can be found in Appendix D.
@@ 7.1 Comparison with CEOs
As demonstrated in Sec. 6.1, the results of CEOs are directly used to accelerate other similarity search processes (Pham & Liu, 2022; Xu & Pham, 2024). In this context, we focus solely on the improvement of CEOs itself. We show that KS1, equipped with the structures $S_{\text{sym}}(m,1)$ and $S_{\text{pol}}(m,1)$, can slightly outperform CEOs($m$) on the original task of CEOs, that is, $k$-MIPS, where $m$ denotes the number of projection vectors and was set to 2048, following the standard configuration of the original CEOs. Since the only difference among the compared approaches is the configuration of the projection vectors, we use a unified algorithm (see construction of projection structure Alg. 4 and MIPS query processing Alg. 5 in Appendix) with the configuration of projection vectors as an input to compare their recall rates. From the results in Table. 1, we observe that: (1) in most cases, KS1 with the two proposed structures achieves slightly better performance than CEOs, supporting our claim that a smaller reference angle yields a more accurate estimation, and (2) $S_{\text{pol}}$ generally achieves a higher recall rate than $S_{\text{sym}}$, verifying that a configuration closer to the best covering yields better performance.
Table 1: Comparison of recall rates (%) for $k$-MIPS, $k$ = 10. The number of projection vectors is 2048. Top-5 projection vectors are probed. Probe@$n$ means top-$n$ points were probed on each probed projection vector. Results are averaged over 10 runs to reduce the bias introduced by random projection.
@@ 7.2 ANNS Performance
We chose ScaNN (Guo et al., 2020), HNSW (Malkov & Yashunin, 2020), and HNSW+PEOs (Lu et al., 2024) as baselines, where ScaNN is a state-of-the-art quantization-based approach that performs better than IVFPQFS, and HNSW+PEOs (Lu et al., 2024) is a state-of-the-art graph-based approach that outperforms FINGER (Chen et al., 2023) and Glass (Zilliz, 2023). Like HNSW+PEOs, KS2 is implemented on HNSW, dubbed HNSW+KS2. DiskANN (Subramanya et al., 2019) is excluded because it is orthogonal to KS2 and focuses on optimization for external storage. ADSampling (Gao & Long, 2023) is excluded because it is designed for a non-SIMD environment. The parameter settings of all compared approaches and Additional experimental results can be found in Appendix D.
@@
Figure 2: Impact of $L$ (See Appendix D.4 for other datasets). $k$ = 10. The y-axis of the upper figures denotes the additional index cost (%) of HNSW+PEOs compared to the original HNSW.
@@ 8 Conclusions
Acknowledgements
This work is partially supported by JSPS Kakenhi JP23K17456, JP23K24850, JP23K25157, JP23K28096, JP25H01117, JP25K00161, and JP25K21272, and JST CREST JPMJCR22M2.
Appendix A Notations
Table 2 lists the main notations used in this paper.
Table 2: Frequently used notations.
Notation Explanation
$\mathcal{D}$ Dataset
$n$ Size of Dataset
$d$ Data dimension
$\bm{v}$ Data vector in $\mathcal{D}$
$\mathbb{S}^{d-1}$ d1d-1 dimensional sphere
$\bm{q}$ Query
$m$ The number of projection vectors
$\epsilon$ Error rate
$S$ Fixed set of $m$ points on $\mathbb{S}^{d-1}$
$H$ Random rotation matrix
$Z_{S}(\bm{v})$ $\mathop{\rm argmax}_{\bm{u}\in S}\left\langle\bm{u},\bm{v}\right\rangle$
$A_{S}(\bm{v})$ $\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle$
$L$ The number of space partitions
$\bm{e}$ Edge in similarity graph
$\bm{w}$ Neighbor vector of $\bm{v}$, connected by edge $\bm{e}$
$\|\bm{e}\|$ Length of edge $e$
$\tau$ Threshold determined by the result list of graph
$\bm{q_{i}}$, $\bm{e_{i}}$ $i$-th sub-vectors of $\bm{q}$ ($1\leq i\leq L$), $\bm{e}$, respectively
$\bm{u}^{i}_{j}$ $j$-th element of the $i$-th sub-vector of $\bm{u}$, where $1\leq i\leq L$ and $1\leq j\leq m$
$\bm{u^{i}_{e[i]}}$ Reference vector of $\bm{e_{i}}$ among all $\{\bm{u^{i}_{j}}\}$’s ($1\leq j\leq m$)
Appendix B Proof of Lemmas
B.1 Proof of Lemma 4.2
@@ Appendix B Proof of Lemmas
$\mathbb{P}[K^{2}_{S}(\bm{q},\bm{v})\geq\cos\theta]=I_{t^{\prime}}\left(\frac{d-2}{2},\frac{d-2}{2}\right)$ (24)
@@
Instead of $\bm{v}$ and $\bm{u}$, we consider $\bm{v}$ and $\bm{u}$, and construct the simplex based on $\bm{q}$, $\bm{-v}$ and $\bm{-u}$ as in Case 1. Then, with the reverse of sign, we finally obtain an equation similar to Eq. (21), except with $\alpha$ replaced by $\pi-\alpha$. The following analysis is similar to that of Case 1.
@@
B.3 Estimation of $J(S_{\mathrm{sym}}(m,L))$
As discussed in Sec 5, for $S_{\text{sym}}(m,L)$, we can establish a relationship between $J(S_{\text{sym}}(m,L))$ and $(m,L)$ as follows.
Lemma B.1.
Suppose that $d$ is divisible by $L$, and $d=Ld^{\prime}$, where $d^{\prime}\geq 3$. Let $c_{d^{\prime}}=\frac{\Gamma\left(\frac{d^{\prime}}{2}\right)}{\sqrt{\pi}\,\Gamma\left(\frac{d^{\prime}-1}{2}\right)}$, $f(y)=c_{d}^{\prime}(1-y^{2})^{\frac{d^{\prime}-3}{2}}$ and $F(y)=\int_{-1}^{y}f(t)\,dt$. We have
The RHS of Ineq. (29) actually denotes $J(S_{\text{ran}})$, where $S_{\text{ran}}$ is the configuration of purely random projections (see Alg. 3 for more detail). A numerical computation of the RHS of Ineq. (29) is shown in Fig. 4 in Appendix.
@@ B.3 Estimation of $J(S_{\mathrm{sym}}(m,L))$
Then we only need to focus on $J(S_{\text{ran}}(m,L))$ and prove that it is equal to the RHS of Ineq. (29). Let $\bm{v}=[\bm{v_{1}},\bm{v_{2}},\cdots,\bm{v_{L}}]\in\mathbb{R}^{n}$ be a vector drawn randomly from $\mathbb{S}^{d-1}$, where $d$ is assumed to be divisible by $L$ and $d=Ld^{\prime}$. Let $r_{L}(\bm{v})$ be the regularized vector w.r.t. $\bm{v}$. That is,
@@
Next, we consider $\mathbb{E}[T^{\prime}(d^{\prime},L)]$. Because of the rotational symmetry of the sphere, the distribution of the inner product $Z=\langle u,v\rangle$ (for fixed $v$ and uniformly random $u$) depends only on the dimension $d^{\prime}$. It has the following density on [1,1][-1,1]:
@@
Appendix C Supplementary Materials
C.1 Explanation of routing test in Ineq. (9)
According to Ineq. (9), there are three core terms I, II, and III, where $\text{I}=\Sigma^{L}_{i=1}{\bm{{q_{i}}^{\top}u^{i}_{e[i]}}}$, $\text{II}=A_{S}(\bm{v})$, and $\text{III}=\frac{\|\bm{w}\|^{2}/2-\tau-\bm{v}^{\top}\bm{q}}{\|\bm{e}\|}$. Here, I/II is exactly the second designed function $K_{S}^{2}$; the computation of I depends on PQ Jégou et al. (2011). III denotes the threshold for $K^{2}_{S}$ and is explained in the original paper of PEOs (see routing test Eq. (9) in Lu et al. (2024)).
C.2 Illustration of Random Projection Techniques for Angle Estimation
@@
C.3 ANNS and Similarity Graphs
As one of the most fundamental problems, approximate nearest neighbor search (ANNS) has seen a surge of interest in recent years, leading to the development of numerous approaches across different paradigms. These include tree-based methods (Curtin et al., 2014), hashing-based methods (Andoni & Indyk, 2008; Lei et al., 2019; Andoni & Beaglehole, 2022; Andoni et al., 2015; Pham & Liu, 2022), vector quantization (VQ)-based methods (Jégou et al., 2011; Ge et al., 2014; Babenko & Lempitsky, 2016; Guo et al., 2020), learning-based methods (Gupta et al., 2022; Li et al., 2023), and graph-based methods (Malkov & Yashunin, 2020; Subramanya et al., 2019; Fu et al., 2019; 2022; Gao & Long, 2023; Peng et al., 2023; Xie et al., 2025; Yang et al., 2025).
Among these, graph-based methods are widely regarded as the state-of-the-art. Currently, three main types of optimizations are used to enhance their search performance: (1) improved edge-selection strategies, (2) more effective routing techniques, and (3) quantization of raw vectors. These approaches are generally orthogonal to one another. Since the proposed structure, PEOs, belongs to the second category, we briefly introduce several highly relevant works below. TOGG-KMC (Xu et al., 2021) and HCNNG (Muñoz et al., 2019) use KD-trees to determine the direction of the query and restrict the search to points in that direction. While this estimation is computationally efficient, it results in relatively low query accuracy, limiting their improvements over HNSW (Malkov & Yashunin, 2020). FINGER (Chen et al., 2023) examines all neighbors and estimates their distances to the query. For each node, FINGER locally generates promising projection vectors to define a subspace, then uses collision counting which is similar to SimHash to approximate distances within each visited subspace. Learn-to-Route (Baranchuk et al., 2019) learns a routing function using auxiliary representations that guide optimal navigation from the starting vertex to the nearest neighbor. Recently, Lu et al. (2024) proposed PEOs, which leverages space partitioning and random projection techniques to estimate a random variable representing the angle between each neighbor and the query vector. By aggregating projection information from multiple subspaces, PEOs substantially reduces the variance of this estimated distribution, significantly improving query accuracy.
C.4 Algorithms Related to KS1 and KS2